### v3

# Can AI writing be salvaged? Mitigating Idiosyncrasies and Improving Human-AI Alignment in the Writing Process through Edits

Abstract.
LLM-based applications are helping people write, and LLM-generated text is making its way into social media, journalism, and our classrooms. However, the differences between LLM-generated and human-written text remain unclear. To explore this, we hired professional writers to edit paragraphs in several creative domains. We first found these writers agree on undesirable idiosyncrasies in LLM-generated text, formalizing it into a seven-category taxonomy (e.g. clichés, unnecessary exposition). Second, we curated the LAMP corpus: 1,057 LLM-generated paragraphs edited by professional writers according to our taxonomy. Analysis of LAMP reveals that none of the LLMs used in our study (GPT4o, Claude-3.5-Sonnet, Llama-3.1-70b) outperform each other in terms of writing quality, revealing common limitations across model families. Third, we explored automatic editing methods to improve LLM-generated text. A large-scale preference annotation confirms that although experts largely prefer text edited by other experts, automatic editing methods show promise in improving alignment between LLM-generated and human-written text.
Human-AI collaboration, Large Language Models, Design Methods, Text Editing, Natural Language Generation, Evaluation, Writing Assistance, Generative AI, Homogenization, Alignment, Behavioral Science
ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Computing methodologies Natural language generation
1. Introduction
Artificial Intelligence (AI) has the potential to revolutionize how we write, communicate, and express ideas (Li et al., 2024). Recent studies have demonstrated the potential of large language models (LLMs) in assisting with various writing tasks, including argumentative (Lee et al., 2022; Dhillon et al., 2024), scientific (Gero et al., 2022), and creative writing (Chakrabarty et al., 2024b; Ippolito et al., 2022b; Yuan et al., 2022; Mirowski et al., 2023, 2024). Aligning LLMs with human preferences (Ouyang et al., 2022) has enabled their transformation into user-friendly tools for non-technical users, such as Google’s WorkSpace Labs, Grammarly, and Sudowrite. However, to truly benefit society, AI writing assistants must enhance human creativity and expression rather than homogenize content or diminish linguistic diversity (Gabriel et al., 2024; Jakesch et al., 2023)
Figure 1. To align models to human preferences, human annotators are typically shown two responses and asked to choose the one they prefer. (i) The top portion of the Figure shows Traditional Alignment: it is often hard to compare two responses that differ widely. (ii) The bottom portion of the Figure shows Alignment via Edits where the original response is edited, allowing for a more granular comparison, with the edited version of the text naturally preferred over the original response.
While LLM-based writing assistants have the potential to improve writing quality and increase author productivity, they also introduce an algorithmic monoculture (Kleinberg and Raghavan, 2021). Padmakumar and He (2023a) and Anderson et al. (2024a) discuss how writing with LLMs unintentionally reduces content diversity, leading to homogenization. This homogenization occurs not just at the semantic level but also at syntactic (structural), lexical (Kobak et al., 2024), and stylistic levels (Shaib et al., 2024). For instance, prior work from Ippolito et al. (2022b); Chakrabarty et al. (2024b) has shown how LLM-generated text is often hackneyed and rife with clichés, while failing to demonstrate rhetorical complexity and often revealing the subtext —a phenomenon known as “telling instead of showing” (Chakrabarty et al., 2024a). Additionally, LLM-generated texts are typically full of redundant exposition, overwrought metaphors, and florid descriptions due to verbosity bias during preference labeling (Saito et al., 2023).
Current AI-assisted writing tools are powered by pre-trained language models that are refined through Reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019). RLHF transforms human preferences into training data to guide language models toward desired outcomes. The most common type of feedback used with RLHF is binary preferences between pairs of examples sampled from one or more language models (Christiano et al., 2017) (See Fig 1 Traditional Alignment). However, this approach has a drawback. The paired outputs may differ in numerous ways and could be equally flawed in containing idiosyncrasies. Asking an annotator to choose between two undesirable outputs does not improve alignment [^1] In the current design of RLHF annotators are not allowed to not pick either. (Casper et al., 2023; Ghosh et al., 2024). We argue that alignment training needs to be aware of how desirable any individual response is, regardless of its preference relationship. Editing undesirable portions of a response can be seen as an effective mechanism for enhancing alignment (See Fig 1 Alignment via Edits). An LLM-generated response that has been edited typically contains fewer undesirable traits and can be paired with the original LLM-generated response for preference ranking ($edited>original$). The challenge however lies in consistently identifying and implementing edits that enhance quality while aligning with human preferences for effective writing. Prompting techniques that encourage the model to self-edit (Madaan et al., 2024) have shown promise however they do not work for long form or paragraph level writing (Pan et al., 2024). The primary reason for this can be accounted to the fact that LLMs don’t inherently know what aspects of the writing needs improvement, how extensively they should edit, or how to make changes that align with any given writer’s expertise level and style.
To address these problems, we first create a comprehensive taxonomy of edit categories based on expert writing practices. We then recruit 18 writers to edit LLM-generated text using categories from our taxonomy. We define edits as changes that alter, replace, or refine specific phrases, clauses, or sentences within a larger text. We restrict our focus to generating text in literary fiction and creative non-fiction, as these genres challenge LLMs with their creativity, emotional nuance, and sophisticated language use. We focus on paragraph-level edits as they balance granularity and scope, reducing costs and annotator fatigue. Paragraphs capture style and context better than sentences, enabling more cohesive improvements. Given LLMs’ limitations in long-term discourse coherence, paragraph-level enhancements facilitate human-AI collaboration. This approach allows humans to guide the overall structure and flow while AI handles lower-level details (Dhillon et al., 2024). Finally, while expert writer’s edits are valuable for identifying idiosyncrasies, this approach may not scale for large text volumes. To address this, we design few-shot prompts that use writer’s edits to identify problematic spans in LLM-generated paragraphs and suggest improvements. This method aims to enhance overall paragraph quality at scale. To summarize our contributions:
  • We propose a comprehensive edit taxonomy informed by expert writing practices that serve as a useful framework to identify and mitigate the idiosyncrasies in LLM-generated text.
  • We release the LAMP (Language model Authored, Manually Polished) corpus containing 1057 ¡instructions,
    response¿ pairs grounded in real-world writing tasks such as Literary Fiction or Creative Non-Fiction. These responses originally generated by LLMs are further edited by 18 professional writers using the above-mentioned taxonomy, resulting in 8035 fine-grained edits (Section 4).
  • We present a thorough analysis of the process of editing LLM-generated text, offering insights into how expert writers edit them, to what extent the edits differ in quantity, how the distribution of edit categories varies across text generated by different model families, and whether LLM generated text contain any specific stylistic idiosyncrasies (Sections 4 and 5.3).
  • We conduct an empirical investigation that tests if LLMs can automatically detect and rewrite their own idiosyncrasies. Our statistically significant results show an encouraging preference trend $Writer-edited>LLM-edited>LLM-generated$ suggesting that edits improve human-AI alignment in the writing process.
  • Finally, we discuss how LLM edits can both mimic and differ from edits provided by professional writers, and what future LLM-based writing support tools can do to improve the co-writing experience.
Our code, data and experimental setup can be found at  [^2] https://github.com/salesforce/creativity_eval/.
2. RELATED WORK
2.1. Text Editing in HCI
Text editing is the process of modifying written content using specialized software. HCI research on text editing aims to improve digital writing tools’ efficiency and usability. Word processors have long allowed flexible editing functions (Pea and Kurland, 1987). Systems like Soylent (Bernstein et al., 2015), MicroWriter (Teevan et al., 2016), and WearWrite (Nebeling et al., 2016) developed interfaces for crowd-based editing, focusing on task breakdown, cost management, and minimizing delays. Robertson and Black (1983) proposes a goal-fate analysis model for text editing behavior, supported by data showing distinct plan units in editing tasks, with potential for intention-based user assistance. Tyler et al. (1982) investigated text editing skill progression and effective training methods. Rosson (1983) examined the impact of experience on editing behavior, questioning if users naturally develop optimal strategies or plateau, noting that experienced users tend to develop more efficient editing heuristics than novices. Reza et al. (2024) present ABScribe, a novel interface that streamlines the process of generating and comparing writing variations using Large Language Models, addressing challenges in existing text-editing workflows and improving writer’s efficiency and satisfaction. Zhou and Sterman (2024) suggest that imperfect AI text suggestions can promote deeper engagement in rewriting, potentially preserving the writer’s authenticity and creative ownership. Park and Lee (2023) found that providing rationales for edits in collaborative writing was generally preferred by participants, despite no significant differences in survey results. This led to design recommendations for effective collaboration. Dang et al. (2022) propose a text editor with continuously updated paragraph-wise summaries as margin annotations to help users plan, structure, and reflect on their writing process.Laban et al. (2023a) introduce InkSync, an LLM-based editing interface suggesting executable document edits. It uses a three-stage approach (Warn, Verify, Audit) to reduce factual errors and enhance editing accuracy, efficiency, and user experience compared to standard chat interfaces. In contrast to existing research, we focus on edits as a method to improve human-AI alignment in writing assistance. Our work characterizes the undesirable aspects of AI writing informed by expert consensus and designs an approach to mitigate these through text editing.
2.2. Text Editing in NLP
NLP research has explored various text editing tasks (Laban et al., 2023b; Pryzant et al., 2020; Choi et al., 2021). Adding to it the advent of Large Language Models has enabled AI-assisted writing tools (Buschek et al., 2021; Ippolito et al., 2022a). Faltings et al. (2021) release the WikiDocEdits dataset and propose an interactive text generation setting in which a user interacts with the system by issuing commands to edit existing text. Raheja et al. (2023) proposed an instruction-based editing system using fine-tuned language models. Shu et al. (2024) developed strategies for cross-sentence rewriting and introduced the OpenRewriteEval benchmark. Reid and Neubig (2022) modeled multi-step editing processes to better mimic human content creation and improve performance on various tasks. Kim et al. (2022) presented a system that iteratively improves fluency, clarity, coherence, and style by detecting editable spans and their corresponding edit intents, then instructing a revision model to refine the text. Yang et al. (2017) developed a taxonomy and classifier for Wikipedia edit intentions. Following them Du et al. (2022) created a multi-domain corpus of revised text with annotated edit intentions. Unlike existing work, we create a resource for text editing that caters to challenging writing tasks (literary fiction and creative nonfiction). Our data consists of 8035 fine-grained edits that are annotated by creative writing experts and we further show how recent advances in few-shot learning can help models improve their own writing by learning from edits provided by the writers.
2.3. Issues in AI Writing
Prior work has highlighted several issues in AI-generated text. Chakrabarty et al. (2024a, b); Ippolito et al. (2022b); Mirowski et al. (2023); Marco et al. (2024) show how LLM-generated text is often hackneyed and rife with clichés, lacks nuance, subtext, and rhetorical complexity. Recent work from Mirowski et al. (2024) shows LLMs fail to act as good creativity support tools for comedy writing and mostly resort to producing bland and biased comedy tropes. They further highlight how existing moderation strategies used in safety filtering and instruction-tuned LLMs reinforce hegemonic viewpoints by erasing minority groups and their perspectives in writing. In summarizing short stories Subbiah et al. (2024) demonstrate how LLMs struggle with specificity and interpretation of difficult subtext. In a similar vein, Tian et al. (2024) shed light on how LLM-generated stories are homogeneously positive and lack tension. Compared to existing work we create a fine-grained taxonomy highlighting the issues in AI writing and further create a large-scale corpus to fuel research in this direction. We also develop automated methods to identify and mitigate issues in AI writing at scale.
2.4. Human AI alignment in Writing
Lee et al. (2024) highlight how AI tools have transformed writing processes, establishing new criteria for future AI writing assistants. In a similar vein Li et al. (2024) reveal that while users benefit from AI assistance in productivity and confidence, potential drawbacks include decreased accountability and diversity in writing. LLMs used in writing assistance can significantly influence human-authored content. Hohenstein and Jung (2020) found LLM-generated text suggestions can affect human writer’s emotional tone. Arnold et al. (2020) showed predictive text encourages predictable writing. Anderson et al. (2024b) and Laban et al. (2023a) found LLMs like ChatGPT helped users generate more detailed ideas, but outputs were less semantically distinct across users (Padmakumar and He, 2023b), and participants felt less responsible for their produced ideas. Recent work from Pan et al. (2024) demonstrates language models can enhance outputs via feedback. However, methods like Iterative Self-refinement scenarios, using another language model as an evaluator, may result in reward hacking, where the model exploits the evaluator’s flaws. For alignment training, it’s crucial to consider the absolute desirability of each potential response, not just how responses compare to one another in terms of preference. Towards this in our work, we create pairs consisting of an initial LLM-generated response and its refined counterparts that by nature is more contrastive (or closely comparable). Our results show that such a pairing results in improved alignment and agreement during preference ranking.
3. DESIGN CONSIDERATIONS TO IMPROVE AI WRITING
The secret to good writing is good editing (to Edit Your Own Writing, 2024). It’s what separates hastily written, randomly punctuated, incoherent rants from learned polemics and op-eds, and cringe-worthy fan fiction from a critically acclaimed novel (to Edit Your Own Writing, 2024). In this section we outline the design principles and desiderata that guided our approach in improving AI writing through textual edits.
Design Principle 1: Develop a comprehensive edit taxonomy grounded in expert writing practices
This principle emphasizes creating a comprehensive taxonomy of edit categories (Faigley and Witte, 1981) rooted in expert editor and writer’s practices. Prior work has shown that experts and novices define revising in very different ways with experts attending more systematically to different aspects of the text than novices (Bracewell et al., 1978; Sommers, 1980; Tyler et al., 1982; Rosson, 1983). By developing such a taxonomy, we aim to provide an approach to analyzing and enhancing LLM-generated text. It also allows for a more granular understanding of the specific areas where AI writing may fall short and enables targeted improvements. Sommers (1980) found that “experienced writers have a second objective; a concern for their readership”. Grounding the taxonomy in expert writing practices ensures that the edits align with the standards of high-quality writing and is acceptable to its readers. Finally this principle also acknowledges the complexity of the editing process, recognizing that different categories of edits may be required at various levels of the text, from sentence-level corrections to broader structural changes (Hayes et al., 1987; Sommers, 1980; Yang et al., 2017).
Design Principle 2: Leverage edits to balance both meaning preservation and substantive semantic changes
Preserving core meaning and intent of the original text is crucial to maintain coherence and faithfulness to the initial ideas. On the other hand, introducing substantive semantic changes is often required to adhere to the quality and characteristics of good writing. Prior work on edit taxonomies focuses on low-level syntactic operations (Faigley and Witte, 1981) or semantic edits (Yang et al., 2017; Daxenberger and Gurevych, 2013; Laban et al., 2023b) tailored to specific websites like Wikipedia. LLM-generated text often benefits from syntactic edits. These edits (primarily meaning preserving) enhance readability by diversifying sentence structures, expanding vocabulary choices, and minimizing repetitive phrasing. Consequently semantic edits (both meaning preserving and changing) in AI writing are important for enhancing specificity or reducing unnecessary flourishes and clichés that can otherwise obscure meaning. Our methodology aims to navigate the tension between maintaining original meaning and introducing necessary improvements to mitigate AI-specific writing quirks.
Design Principle 3: Utilize edits as a mechanism for enhancing human-AI alignment in writing
Current AI writing systems are developed using pre-trained language models (LMs) refined through human interaction, employing supervised learning and reinforcement learning (RL) techniques. Reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019) is a key approach, transforming human input into training data to guide LMs toward desired outcomes. The most common type of feedback used with RLHF is binary preferences between pairs of examples sampled from one or more Language Models (Christiano et al., 2017). However, a learned preference ordering can fail to converge to the true one when the desirability of examples depends on noise (Gao et al., 2024). To tackle this we propose edits as a mechanism for enhancing alignment. An LLM-generated response that has been edited typically contains fewer undesirable traits and can be paired with the original LLM-generated response for preference ranking ($edited>original$). This method of incorporating edits for improved preference data collection has also been adopted by contemporaneous work from Meta AI (Dubey et al., 2024) for training Llama3.1 as well as Contextual AI (D’Oosterlinck et al., 2024).
4. LARGE SCALE DATA COLLECTION PROCESS
Figure 2. The pipeline for data creation. Step 1) Extracting context-independent paragraphs from our respective sources Step 2) Using an LLM to automatically generate instructions for corresponding human-written text Step 3) Use the generated instructions grounded in real-world writing to elicit responses from LLMs to create $<$instructions,response$>$ pairs
We aim to create a valid test-bed to evaluate the quality of LLM-generated text on realistic writing tasks that require creative skill. We follow a three-step approach illustrated in Figure 2: (1) First we select original paragraphs of human-written text from trusted publication venues, (2) Second we reverse-engineer each of these paragraphs into a writing instruction. Because each instruction originates from an existing human-written paragraph within a piece of creative writing, this simulates real-world writing situations. (3) Third, we prompt several LLMs to generate responses to each of the writing instructions. In the following subsections, we first detail each of these steps, and then describe the formative study we conducted to develop a taxonomy of idiosyncrasies in LLM-generated text.
4.1. Collecting Instruction and Response Pairs
Domain Source & Genre Example Seed Paragraph with Generated Instruction #N
Fiction The NewYorker (Literary Fiction)
The sunset is a red-gold rumpus on the western sky. It has rained. The crow tosses
itself from branch to branch, pole to pole, glistening on its pace, and she follows. They
are soon far from where they began, streets unfamiliar to her, an older part of town,[…..
….]. A man reading. Old Christmas tree in a corner. It feels secret. The sky is
clearing overhead. She feels secret, too. She feels tremendous.
Instruction : Can you describe a vivid scene at sunset that transitions into
nighttime, incorporating elements of nature, urban surroundings, and
personal observations?
815
Creative Non-Fiction NYTimes (Travel Writing)
Prague, the Czech capital, is finding a new balance between preserving its past and
embracing the future, improving many of its important historic sites while making
striking additions to its skyline. [………] Stop by for a coffee, hit up one of the many
great new bakeries or visit a charismatic old beer hall as you explore a city that is
clearly entering its prime.
Instruction : How is Prague balancing its historical preservation with modern
development while enhancing local amenities and vibrant
neighborhoods outside Old Town?
110
NYTimes (Food Writing)
The origins of the fruit sandwich are believed to go back to Japan’s luxury fruit stores &
the fruit parlors attached to them. This version comes from Yudai Kanayama, a native of
Hokkaido who runs the restaurants the Izakaya NYC and Dr Clark in New York. [………]
.The sandwich looks like dessert but isn’t, or not exactly; it makes for a lovely little meal
that feels slightly illicit, as if for a moment there are no rules
Instruction : How did Yudai Kanayama reinvent the traditional Japanese fruit
sandwich to create a unique culinary experience?
83
NYTimes (Personal Essay)
My dad’s Hinge profile showed his pandemic scruff, cheeky smirk and favorite frayed
T-shirt. He claimed his strength was listening [………] We revamped his prompts to
highlight his superhero dad qualities and love of movies. My dad’s Hinge profile no
longer seemed unhinged. Two months later, he had a girlfriend.
Instruction : How did the changes to your dad’s Hinge profile, including updated
photos and revamped prompts, impact his success on the dating app?
19
Dear Sugar (Internet Advice)
What is a prestigious college? What did attending such a school allow you to believe
about yourself? What assumptions do you have about the colleges that you would not
describe as prestigious? What sorts of people go to prestigious colleges and not [………]
I believe our early experiences and beliefs about our place in the world inform who we
think we are and what we deserve and by what means it should be given to us.
Instruction : How do your beliefs and assumptions about educational privilege
and the type of schools people attend shape your current view of yourself
and others?
30
Table 1. Venues of source material used to extract real-world examples of creative writing, along with example seed paragraphs and the generated instructions and the number of samples per genre.
To curate source material, we select five well-regarded  [^3] https://www.pewresearch.org/politics/2012/09/27/section-4-demographics-and-political-views-of-news-audiences/ publication venues, listed in Table 1, that publish pieces in different domains ranging from fiction to food writing and internet advice. For each venue, we manually extract between 100-700 pieces of writing and isolate individual paragraphs. We then manually review these paragraphs, ensuring they are long enough and can stand alone as coherent pieces of writing without requiring additional context. In total, we selected approximately 1200 paragraphs following this procedure. The Literary Fiction genre has a larger representation (80%) in our selection, while the creative non-fiction genres have a smaller representation.
Next, we follow Li et al. (2023)’s approach of Instruction Backtranslation to automatically generate instructions corresponding to each of the selected paragraphs. Specifically, we prompt an LLM (i.e., GPT4o) to summarize each paragraph into an open-ended question. Questions obtained through back-translation (see examples in Table 1) can be interpreted as realistic writing instructions. We manually verified the generated instructions, filtering out questions that were ill-formed or overly specific, yielding a total of 1,057 writing instructions.
Finally, we use the generated instructions to elicit responses from three state-of-the-art LLMs: OpenAI’s GPT-4o (OpenAI, 2024), Anthropic’s Claude-3.5-Sonnet (Anthropic, 2024), and Meta’s Llama 3.1-70b (Dubey et al., 2024). Each LLM is used to generate responses to one-third of our instruction data. We ensure that each LLM responds to instruction across all domains in equal proportion. To generate high-quality responses, we provide each LLM with the writing instruction, as well as the genre and source, and instruct it to adhere to the style of the venue. The prompt further specifies: “Try your best to be original, avoiding clichés or overused tropes. Do not use ornamental language and focus on nuance, simplicity, and subtext” (See Prompt in Appendix A Table 13). Through this process we obtain 1,057 writing $<$instructions, response$>$ pairs, with responses averaging 205 words. This collection of instructions and LLM-generated responses serves as the foundation for the three studies we conducted: the formative study, the full-scale editing annotation, and the preference annotation.
4.2. Formative study: formulating the taxonomy for fine-grained edits
Our formative study observed writers with copy-editing experience as they edited LLM-generated text in the creative writing domain. We aimed to identify common edit categories. The study consisted of three phases. First, participants were individually briefed via video conference on the study’s objectives. Next, they accessed a web application (Figure 9) to view $<$instruction, response$>$ pairs from our dataset (Section 5). For each sample, participants highlighted problematic response spans, suggested rewrites, and tagged each span with a free-form category to characterize the issue. We recruited eight participants for the formative study, with each completing annotations for 25 samples.
ID Profession Gender Age Educational Background
W1 Writer & Editor at Magazine Male 28 MFA in Fiction
W2 Writer & Fellow at Fine Arts Work Center Male 29 MFA in Fiction
W3 MFA Fiction Student Male 31 MFA in Fiction
W4 Writer Female 30 MFA in Non-Fiction
W5 Writer Female 35 MFA in Fiction
W6 MFA Poetry Student Female 27 MFA in Poetry
W7 Writer & Journalist Female 35 MFA in Non-Fiction
W8 MFA Fiction Student Male 26 MFA in Fiction
Table 2. Pilot study: background of participants.
For participant selection in our formative study, we limited involvement to individuals with established expertise in creative writing. Participants were required to have completed a Master of Fine Arts (MFA) in Creative Writing and were recruited through mailing lists from MFA writing programs in the United States. This aligns with prior work from Chakrabarty et al. (2024a), using the Consensual Assessment Technique (Amabile, 1982), which emphasizes the importance of recruiting domain experts. During the initial video call, we confirmed participants’ familiarity with copy-editing and informed them they would edit LLM-generated texts. We recruited participants through UserInterviews  [^4] https://www.userinterviews.com, a professional freelancing platform, paying \$75 USD for study completion. Editing a response took 4-6 minutes, with all participants finishing within two hours. Table 2 shows diverse demographics and professional backgrounds of the recruited creative writers. [^5] The research was conducted at an institution without a formal IRB approval process. However, an Ethical Practices team reviewed the work and study protocols. No personally identifiable information (PII) was collected or shared during data collection, and participants were offered compensation regardless of study completion. In total, the eight participants edited 200 paragraphs, annotating roughly 1,600 edits attributed to 50 distinct initial categories. We used this data as the foundation for our next analysis, which aimed to develop a taxonomy for categorizing edits.
4.3. From initial to final categorization of edits
Final Category Initial Categories Participants
Cliche Cliched image from old westerns, Cliche, Hackneyed
W1,W2,W3,W4,
W5,W6,W7,W8
Unnecessary/Redundant
Exposition
Repetition of what has already been stated, Unnecessary, Show don’t tell,
Repetition, Cut Unnecessary, Unnecessary because implied, Over
exposition, Fluff, Slim down, Trying to cut things down, Concision
W1,W2,W3,W4,
W5,W6,W7,W8
Purple Prose
Too wordy, Purple Prose, Ornamental, Very Verbose, Clunky
Unnecessarily wordy, Simplify, Overwrought, Mixed metaphor
W1,W3, W4
W5,W7,W8
Poor Sentence Structure
Structure, Transition, Editing for clarity, Better to split up into two
sentences, Run-on sentence, Very Long and Complex Sentence
W1, W5, W7,W8
Lack of Specificity
and Detail
Lacks specificity, Overly General, More details to help move the reader,
Added details, Creating a scene, Contextualizing information,
Deepening internality, Needs to be more specific, Adding Voice
W1,W2,W3,W4
W5,W6,W8
Awkward Word Choice
and Phrasing
Word Choice, Pronoun Clarity, Passive, Awkward Word Choice,
Wrong choice of word, Rewording, Rephrasing, Weird Phrasing,
Inelegant
W1,W2,W3,W4
W5,W6,W7,W8
Tense Inconsistency
Fragment sounds weird-is it past or present?, Wrong Tense,
Inconsistent Tense
W1, W5, W7,W8
Table 3. Our final taxonomy for fine-grained edits to mitigate idiosyncrasies in AI writing
We observed significant semantic overlap among the 50 initial categories used by participants, suggesting potential consolidation into a unified taxonomy. For instance, “Show don’t tell” (W4) corresponded to “Unnecessary because implied” (W6). Using a general inductive approach for qualitative data analysis (Thomas, 2006), we synthesized these 50 initial categories into a comprehensive, fine-grained taxonomy of edits. First, two authors independently bucketed these categories into initial low-level groups. Through iterative discussions, these groups were refined to reduce overlap and establish shared groupings. The refined low-level groups were then aggregated into high-level categories. Each high-level category was assigned a name reflecting its generalized representation.
The aggregation process yielded 7 distinct edit categories, presented in Table 3 along with contributing participant IDs. Final categories were retained only if derived from initial categories identified by at least four participants, ensuring majority representation in editing feedback  [^6] Only 5% of edit categories were not included in the 7 categories as they did not have enough coverage. It’s worth noting that not every LLM-generated response exhibits all these idiosyncrasies. The formative study’s objective was not to establish the relative prevalence of each category. Instead, this taxonomy serves as a useful framework when considering the categories of edits to apply to LLM-generated content. The categorization provides a structured approach to refining such text.
4.4. Final Taxonomy for Fine-Grained Edits
Here we describe our final taxonomy for fine-grained edits. Table 5 shows examples of edits in each of these categories defined below
4.4.1. Cliché
Clichés in writing are pejoratively characterized as phrases, ideas, or sentences overused to the point of losing their original impact or meaning. They often use vivid analogies or exaggerations from everyday life to describe abstract concepts. While occasionally effective when used sparingly, the frequent use of clichés in writing is generally viewed as a sign of inexperience or lack of originality (Fountain, 2012). Replacing clichés with fresh, original language improves the writing and engages readers more effectively.
4.4.2. Unnecessary/Redundant Exposition
Unnecessary or redundant exposition refers to the inclusion of excessive, repetitive, or implied information in writing. This common pitfall often involves restating the obvious or providing details that add little value. In a conversation with W2 they said “I’m adding a category of edit called “fluff” - this is a common term in the writing world to refer to unnecessary filler”. Effective writing embraces the principle of “show, don’t tell,” allowing readers to infer meaning from context rather than relying on explicit explanations (Noble, 2013; Chekhov, 1984; King, 2000; Burroway et al., 2019). Impactful writing, often allows the core message to shine through without being obscured by unnecessary verbiage.
4.4.3. Purple Prose
In literary criticism, purple prose refers to excessively elaborate writing that disrupts the narrative flow by attracting undue attention to its flamboyant style (wik, 2018). This can detract from the text’s overall appreciation. Such writing is often difficult to read, using sprawling sentences, abstract words, and excessive adjectives, adverbs, and metaphors to convey little information. Careful editing can trim purple prose by replacing ornate language with more direct expressions, resulting in clearer writing that preserves narrative flow and the author’s voice.
4.4.4. Poor Sentence Structure
Poor sentence structure reduces the clarity and readability of writing (Bormuth, 1966; Klare, 1974; Meyer, 2003). A lack of proper transitions can make the text feel disjointed and hard to follow. Editing for clarity (Dermer et al., 2009) often reveals that it’s better to split a convoluted thought into two sentences, rather than forcing it into one (Lane and Lange, 1999). Run-on sentences, characterized by multiple independent clauses improperly connected, are also frequent problems in AI writing (Chakrabarty et al., 2024a). These, very long and complex sentences can overwhelm the reader, making the core message difficult to grasp. Edits that reduce these problems lead to more coherent and fluent text.
4.4.5. Lack of Specificity and Detail
Lack of specificity and details in writing often stems from a writer’s tendency to rely on broad generalizations (MacDonald, 1986). This overly general approach fails to engage readers, leaving them unable to visualize scenes or connect with any given writing on a deeper level. Good writing often focuses on adding vivid details that create a clear mental image (Doty, 2014; Fleckenstein, 1991; Kuzmičová, 2014), contextualizing information to give it relevance (Rababah, 2022), and deepening the internality of characters or subjects (Currie, 1990; Fishelov, 1990). Additionally, developing a unique voice through carefully chosen words and phrases can inject personality into the writing, making it more engaging and distinctive (Nelson and Castelló, 2012; Herring, 2007). Edits belonging to this category typically make the text longer as writers add more details to make the text engaging.
4.4.6. Awkward Word Choice and Phrasing
Awkward phrasing can significantly reduce writing quality, often confusing or disengaging readers. This issue typically involves misused or disproportionate use of certain words (Kobak et al., 2024), unclear pronoun references, or an overuse of passive voice. In an email W1 pointed out “Another little observation to share: a very common phrasing in these excerpts is ‘seem to _(verb)_’. This is not technically wrong it’s just inelegant, something many writing teachers have told me to avoid. Unless there is some specific uncertainty or doubt about the verb action, it’s always preferred to just use the verb without ‘seem’ (ex. from the current excerpt I have up: ‘amplified’ is better than ‘seemed to amplify’)”. Editing plays a crucial role in refining these elements. Through careful revision, writers can identify and replace imprecise or ill-fitting words with more appropriate alternatives, ensuring each term accurately conveys the intended meaning.
4.4.7. Tense Inconsistency
Tense inconsistency is a prevalent issue in writing. It occurs when a writer inadvertently shifts between past, present, and future tenses often even within the same paragraph or sentence. This grammatical misstep can make the timeline of events unclear and detract from the overall coherence of the text. Careful editing plays a crucial role in addressing this issue. By paying close attention to verb forms and temporal indicators, editors can improve writing that deals with tense inconsistency (Moens, 1987).
[…]
Figure 3. Interface to collect edits from writers on LLM-generated text
ID Profession Gender Age Educational Background Responses_Edited
W1 Writer & Editor at Magazine Male 28 MFA in Fiction 123
W2 Writer & Fellow at Fine Arts Work Center Male 29 MFA in Fiction 109
W3 Writer & Teacher Male 31 MFA in Fiction 119
W4 MFA Fiction Student & Translator Male 30 MFA in Fiction 25
W5 Writer Female 35 MFA in Poetry 77
W6 MFA Poetry Student Female 27 MFA in Poetry 23
W7 Writer & Journalist & MFA Fiction Student Female 35 MFA in Fiction 71
W8 MFA Fiction Student Male 26 MFA in Fiction 23
W9 Writer & Editor Male 30 MFA in Fiction 25
W10 Writer & Creative Writing Instructor Female 28 MFA in Fiction 25
W11 Writer Female 27 MFA in Poetry 25
W12 Writer & High School Teacher Male 33 MFA in Fiction 24
W13 Writer & Editor Female 29 MFA in Non-Fiction 25
W14 MFA Fiction Student Male 26 MFA in Fiction 24
W15 Poet Female 28 MFA in Poetry 125
W16 Writer & Director Male 31 MFA in Literary Arts 122
W17 Writer Non-Binary 28 MFA in Poetry 25
W18 Screenwriter & MFA Literary Arts Student Female 27 MFA in Literary Arts 67
Table 4. Background of participants who provide span level edits on LLM generated responses
4.5. Collecting Edits on LLM Generated responses
Categorization Paragraph with rewrites
Cliche
As Sarah stepped off the bus, the scent of pine and damp earth enveloped her. [……] In the kitchen, she found
herself reaching for the cabinet where her mother always kept the coffee, only to stop short. The realization
that she was alone here, truly alone, settled over her like a heavy blanket. This time, though, she was
alone. Her mother would never come back.She sank into a chair at the old oak table[….]
Unnecessary
/ Redundant
Exposition
As Mingus and Dylan stepped out of the car, […] The Brooklyn-Queens Expressway loomed above, a concrete
behemoth that cast long shadows over the desolate landscape. cast a long shadow.[…] For a moment, he
stood there, lost in thought, as the city seemed to hold its breath around him.
Lack of
Specificity
and Detail
Dr. Arthur Steiger’s fall from grace began with a series of whispered concerns among his colleagues
at Cormac General Hospital. Pain was Dr. Arthur Steiger’s forte. Not inflicting it, that is, but resolving
it. Whenever a patient had problem, whether a tear in a tendon, a sprain, a knock, a headache, a
broken bone – it was Dr. Steiger that knew what to do.The small-town pain specialist had always been
known for his compassionate approach, but as opioid addiction rates climbed in the community [….]
Poor Sentence
Structure
As the night wore on, Z.’s laughter grew louder, his words slurring together like a sloppy melody. N.
and I exchanged a knowing glance, our concern simmering beneath the surface.Z. was drinking more
and more as the night went on. He laughed more loudly. His words started to slur, blurring one into
the next.I looked at N., who knew what I was thinking. We were going to have to take care of him. At
first, it was just a slight stumble, a misstep that could be brushed off as a joke. But as the hours passed,[…]
Purple Prose
My mother cried not just because twenty grand vanished into the ether[…..].All of it vanished, cycling back
through her mind, not as numbers but memories of scraped knees she bandaged alone and birthdays
where her absence was felt more acutely than her presence.The sobs emerged from this deep well of
unspoken expectations, leaving behind a residue of weary resilience and a few hopeful echoes yet
unwilling to completely extinguish. She cried. She cried deep from this well of scraped knees she
bandaged alone and birthdays she missed to work. She cried for unfairness. She cried without relief.
Awkward Word
Choice and
Phrasing
I remember the city as a place of perpetual twilight, where the sky seemed to hover hovered between dawn
and dusk [….] glass towers, and the sound music of sirens […..] bodega on the opposite side still sold reeked
of warm beer and stale cigarettes. The people were a blur of faces, each with their own story of [….]
Tense
Inconsistency
As the sun dipped below the horizon, Elliot found himself engulfed by the growing darkness on Route 7.
The first snowflakes began to drift drifted down from the heavens,[…]
Table 5. Example of Edit types from our data
With the finalized taxonomy of edits, we next conducted a larger-scale annotation study. The purpose of this study was to collect edits from writers on LLM-generated responses, categorizing them according to the established taxonomy.
This task followed a similar format to the formative study where participants were provided access to an editing interface (Figures 3) populated with instructions and LLM-generated responses. In this interface, participants could select any span of text in the response and suggest a rewrite. Unlike the formative study, participants had to choose from the seven predefined categories in our taxonomy for each edit, rather than entering free-text categories. Participants received training about the taxonomy via email before beginning annotation. The training incorporated example edits for each category, akin to those in Table 5. Participants had no set limit on edits per response but were urged to improve the text as they saw fit. The interface logged all edits chronologically and offered an undo feature, enabling us to track the entire editing process, not just the final product.
After completing their edits, participants assigned two scores to the sample: an Initial Writing Quality Score (IWQS) for the original response quality, and a Final Writing Quality Score (FWQS) for the post-edit quality. Both used a 1-10 scale, with 1 being lowest and 10 being the highest quality. The scores were incorporated to add a quantitative dimension to the qualitative process of editing. Additionally, the self-reported writing quality scoring system serves as a signal for writers to recognize their own improvements, set personal goals, and develop intrinsic motivation for enhancing their work.
Editing is a personal, time-consuming task, with edit quality dependent on participants having sufficient time to carefully read and consider improvements. To ensure quality, we maintained communication with all recruited participants. Participants completed the task in batches of 25, which typically took 3 hours, and were compensated \$100 USD for each batch. We recruited 18 writers with formal creative writing backgrounds from MFA mailing lists for our study, including 3 participants from the formative study (Table 4). Over 2.5 months, these writers edited LLM-generated responses based on their availability. Due to staggered start times, the number of edited samples varied among participants (see details in Table 4). In total, each of the 1,057 ¡instruction, response¿ pairs we had prepared was edited by at least one participant, and 50 responses were edited by three participants, allowing us to study similarities and differences that occur when multiple writers edit the same response. The next section details the analysis we performed on the 8,000+ collected edits.
5. The LAMP Corpus
5.1. Overall Statistics
We created the LAMP Corpus by collaborating with 18 writers who edited 1,057 LLM-generated paragraphs, gathering about 8 edits per paragraph, totaling 8,035 fine-grained edits. The data includes paragraphs from Claude3.5 Sonnet (368), GPT4o (393), and Llama3.1-70B (296). Figures 4-5 present analyses of the LAMP Corpus, offering insights into how professional writers edit LLM-generated text and revealing a surprising lack of difference in writing quality across different model families (Zhou et al., 2024).
Figure 4. Analysis of 1,057 paragraphs edited by 18 Writer participants, analyzing: (4(a)) the edit operations they perform (insertions, deletions, etc.), (4(b)) the writing quality scores they assign, (4(c)) comparing writing quality scores across LLMs, (4(d)) the relationship between IWQS and editing amount.
Figure 5. (5(a)) the categories of edits they implement, and (5(b)) the relationship between writing quality scores and error categories.
We analyze the editing process by examining edit operations: insertion, deletion, or replacement. An edit is an insertion if it deletes no characters and adds 40+ characters net. Conversely, it’s a deletion if it adds no characters and removes 40+ characters net. All other edits are replacements. Figure 4(a) shows edit operations by participant for each paragraph. Replacements are most frequent (74%), followed by deletions (18%) and insertions (8%). Editing styles vary: some participants primarily use replacements (W2, W9, W10, W16, W18), while others employ deletions more often (W1, W5, W7, W8, W13, W17). Insertions are uncommon across all participants. To quantify meaning-preserving vs. meaning-changing edits, we calculate semantic similarity between original and edited text using BERT score(Zhang et al., 2019). Using a threshold of 0.6  [^7] This threshold was decided by manually analyzing 100 edits, we classify edits with similarity ¿ 0.6 as meaning-preserving. Of 6468 non-deletion edits, 70% are meaning-preserving, with the rest meaning-changing. This finding supports our Design Principle 2.
Meaning Preserving During the quarantine, the days stretched like endless corridors, each more indistinguishable from the last 0.72
The days blurred into themselves during the quarantine, and I couldn’t tell one from the other
She glanced down the hallway, suddenly aware of how quiet it was for a Tuesday evening 0.79
It was eerily quiet for a Tuesday evening
Meaning Changing
Sophia took her smoking breaks in the back garden, a ritual she kept as precise as the time on
the old clock in her kitchen
0.52
One of the great comforts of old age was the ability to stop caring what other people thought
a brief reprieve from the unsaid words that floated between them 0.48
time to think of what to say. She hadn’t told them her boyfriend was black
Table 6. Examples of Meaning Preserving vs Meaning Changing Edits. Each example is a pair of original and edited span. Last Column shows the semantic similarity (BERT scores)
The annotation interface allowed participants to provide Initial and Final Writing Quality Scores (IWQS and FWQS) for each paragraph, ranging from 1 to 10. Figure 4(b) shows the distribution of these scores for each participant, revealing significant variability (e.g., W1’s median IWQS is 7, W18’s is 2). Calibration of writing quality scores is a known challenge, and we follow prior work in normalizing the scores into z-Scores by subtracting the mean, dividing by the standard deviation of the scores for each participant (Graham et al., 2013; Maddela et al., 2022), and re-scaling them to the 1 to 10 range. Subsequent analyses use these normalized scores.
We compute an edit distance between the original LLM-generated text and the final edited text, by calculating a character-level Levenshtein distance (Levenshtein, 1966) between the two strings of texts. The edit distance measures the “amount of editing work” performed by a writer. Figure 4(d) shows a negative correlation between edit distance and IWQS (Pearson’s $r=-0.31$), indicating that higher perceived text quality (high IWQS) requires less editing, while lower IWQS necessitates more editing.
Figure 4(c) shows the average IWQS for each LLM on creative non-fiction and fiction writing tasks. Writers were unaware of which model generated each text, and tasks were shuffled to avoid bias. This analysis estimates the writing quality of the three models in both domains. Comparing model scores, we find no significant difference in writing quality across the three models. GPT-4o and Claude 3.5-Sonnet perform slightly better on creative non-fiction instructions (average 5.2) compared to Llama3.1-70B (5.0), though the difference is not statistically significant. All models show a slight decrease in performance for fictional instructions, with an average IWQS of 4.5. This suggests fiction writing may be more challenging for LLMs than creative non-fiction. These findings differ from task-oriented benchmarks that reveal performance gaps between models in areas like factual or logical reasoning. Our results indicate that, when it comes to creative writing, writers perceived no significant qualitative differences among the texts generated by large language models (LLMs) such as GPT-4, Claude 3.5 Sonnet, and Llama 3.1 70B.
Figure 5(a) displays edit categories applied by writers to texts from three LLMs. The distribution is similar across models, with the most common categories being Awkward Word Choice and Phrasing (28%), Poor Sentence Structure (20%), Unnecessary/Redundant Exposition (18%), and Clichés (17%). Minor differences include GPT-4o using more purple prose and Llama3.1-70B generating more unnecessary exposition. Overall, LLMs across the three model families exhibit similar idiosyncrasies that are edited out in similar proportions by professional writers. Figure 5(b) illustrates the relationship between edit categories and IWQS. Higher IWQS scores correspond to fewer total edits, with texts rated 2 averaging 10.2 edits and those rated 10 receiving 2.4 edits, confirming that higher-quality texts need less editing. This trend however varies across edit categories: “Unnecessary/Redundant Exposition” and “Lack of Specificity and Detail” remain relatively constant, while the number of “Awkward Word Choice and Phrasing” and “Cliché” edits decrease as IWQS increases, suggesting a stronger correlation with perceived writing quality.
5.2. Writers differ greatly in the amount of editing they do: But to what extent?
W3 W12 W16
Original
,the numbers glaring back at
me like an unsolvable riddle
glaring back at me like an unsolvable riddle
,the numbers glaring back at
me like an unsolvable riddle
Category Cliche Cliche
Unnecessary/Redundant
Exposition
Edited . The numbers stared back
barreling over one another as they raced
to some unseemly height.
Original
and an unsettling sense of
mystery that gnawed at me
more than the inexplicable
weight itself
Her words felt like a placeholder for an
answer neither of us had yet. I walked out
with a slip for blood tests and an unsettling
sense of mystery that gnawed at me more
than the inexplicable weight itself.
Category Purple Prose Cliche
Unnecessary/Redundant
Exposition
Edited -
But when I saw her turn to go, whispering
in the halls with a colleague, I knew there
was still something she had yet to tell me.
Table 7. Original spans selected by 3 writers from the same paragraph. ‘-’ denotes the span was deleted while editing
The writer’s approaches to editing vary based on personal or organizational philosophy. Some prioritize preserving the original voice and make minimal changes to preserve authenticity (Tanselle, 1976). Others may take a more interventionist stance, heavily revising to align with their vision or house style. Additionally, some writers might make fewer but more impactful changes, while some might make numerous small revisions. To quantify this, we asked 3 writers (W3, W12, and W16) to edit a subset of the same 50 paragraphs from the LAMP Corpus. As expected, these three writers differed in the amount of editing they did. W3 did 9.4 edits on average while W12 and W16 did 6.0 and 6.3 edits on average. On average the span level precision (see Section 6.1 for more detail on the metric) between the 3 writers was 0.57 suggesting a moderately significant agreement.
Table 7 shows how sometimes writers select the exact same problematic span but assign different categories. For a span both W3 and W16 selected “,the numbers glaring back at me like an unsolvable riddle” (Table 7 Row 1) but chose separate categories (Table 7 Row 2). Similarly, W3 and W16 both choose the exact same span, but different categorizations. It should however be noted that both categorizations can be correct interpretations. When one relies on overused phrases or clichés, they often state the obvious or provide information that readers can easily infer implicitly. This results in redundant or superfluous exposition that doesn’t add value to the narrative. Other times writers may select the same category but with only partially overlap on the selected span (Table 7 Row 1 W3 vs W12). Looking at (Table 7 Row 4 and 5; W3 vs W12) there is a partial overlap in the selected span “and an unsettling sense of mystery that gnawed at me more than the inexplicable weight itself”. However, the selected categories are Purple Prose and Cliché respectively. Here again, it should be noted that Purple Prose is a style of writing that can be original or cliché, depending on its usage, context, and frequency. Not all elaborate writing is overused, but when certain ornate phrases or styles become too common, they can cross the line into cliché territory. W16 however did not edit this span.
We also highlight that diversity in edits among writers such as selecting different spans or rewriting it in an individualistic style is a positive aspect that prevents homogenization while still improving LLM-generated text as shown by our results in Section 6.2.
5.3. Are there any specific stylistic idiosyncrasies in LLM generated responses?
Syntactic Pattern
% of
Times
Edited
Representative Sequence
DT NN IN NN CC 54%
a mix of pride and, a mix of fear and, a sense of protection and, a sense of wonder and,
a means of connection and, a pang of nostalgia and, a pang of disappointment, but,
a flicker of hope or, a blend of relaxation and, a blend of curiosity and, the power of
storytelling, the power of empathy, and, a web of belonging and [….]
NN IN NN CC NN 35
scene of chaos and destruction, mix of desperation and resolve, mix of relief and gratitude,
torn between curiosity and caution, breakfast of bread and jam, sense of calm and normalcy,
perception of loyalty and identity, glimmer of fear and vulnerability, meaning of protection
and care, story of struggle and resilience, blend of fear and hope, [….]
DT JJ NN IN PRP\$ 40
a constant reminder of his, the mundane routine of our, the intricate tapestry of its,
the subtle shift in their, the potential weight of its, a quiet sigh as her,
a small acknowledgment of their, the upcoming chapter of her, a silent battle between his,
a complex blend of their, the subtle shift in her, the unspoken plea in her […]
DT NN IN JJ NN 27
the fabric of daily life, a moment of genuine connection, a life of absolute relaxation,
the face of inevitable loss, the weight of past grievances, a state of constant unease,
a residue of weary resilience, a sea of unspoken expectations, a mask of controlled concern,
the weight of unresolved history, a foundation of silent understanding, [….]
IN DT NN IN NN CC 45
with a mix of wariness and, by the hum of traffic and, in a flurry of pursuit and,
into a world of precision and, in a gesture of comfort and, in a storm of pain and,
for the sake of stability and, in the rhythm of routine or, in the magic of family and,
like the depth of understanding and, with a sense of nuance and […]
Table 8. Idiosyncratic Sequences following certain syntactic patterns in LLM generated responses that are edited by writers. These syntactic patterns do not occur in the human written seed paragraphs
Recent work from Shaib et al. (2024) uses syntactic patterns with Part-of-speech  [^8] https://www.sketchengine.eu/blog/pos-tags/ as abstract representations of texts, that can capture more subtle repetitions than mere text memorization. They find that language models tend to use repetitive syntactic templates more often than humans and these patterns can help evaluate style memorization in language models. Following their experiments we consider Part-of-speech templates of length $n\epsilon\{5,6,7,8\}$ in LLM-generated responses as well as the original seed human-written paragraphs (Table 1). We looked at the 50 most common templates in LLM-generated responses and found that 15 templates do not occur as frequently in original human-written seed paragraphs. Table 8 shows representative sequences corresponding to particular syntactic patterns present in higher proportion in LLM-generated responses. These sequences constitute categories of Clichés, Unnecessary/Redundant Exposition or Poor Sentence Structure and are often heavily edited by writers in our study.
Figure 6. Distribution of peculiar and odd words and phrases occurring in LLM-generated text vs. human-written text in the LAMP Corpus.
To better understand idiosyncrasies, we examined awkward words/phrases occurring disproportionately in LLM-generated responses. For instance, Figure 6 shows how a word like unspoken occurs in about 15% of LLM-generated responses. Similarly phrases such as weight of, sense of, mix of occur very rarely or not at all in original seed paragraphs (Table 1) while they occur frequently in LLM-generated responses. We also found peculiar and uncommon phrases generated by LLMs across several responses such as air was thick, hung in the air, eyes darting, a sense of unease (grew/growing/settles) in the pit of (her/my) stomach. The most surprising finding is that all 3 LLMs generate these idiosyncratic words/phrases suggesting possible overlap/mixture in instruction tuning data across model families or one model trained on synthetic data generated from another model (Zhou et al., 2024).
6. AUTOMATIC DETECTION AND REWRITING OF LLM IDIOSYNCRASIES
Expert text editing can analyze and reduce LLM idiosyncrasies at a small scale, but automated methods are needed for resolution at a larger scale. Building on Hayes et al. (1987) and Scardamalia (1983), we develop techniques to separate detection and rewriting tasks, evaluating them using LAMP Corpus annotations. Given automated evaluation limitations for text editing(Dou et al., 2023), we conduct a large-scale preference annotation study with LAMP Corpus writers, comparing human and LLM-produced edits. To accommodate methods that require training samples, we split our data: 146 of 1057 LAMP Corpus paragraphs for training, the rest for testing.
6.1. Automatic detection of problematic spans in LLM-generated text
We formulate the problem of detecting problematic spans in LLM-generated text as a multi-span categorical extraction problem. In other words, given a paragraph of LLM-generated text, the method must output a list of non-overlapping spans present in the original text, and assign a category to each extracted span (from the list of categories of the LAMP Corpus).
To evaluate various methods, we use the span-level precision metric, a common metric used in NLP tasks requiring comparison of extracted spans (Rajpurkar et al., 2016). Span-level precision measures the degree of overlap between predicted spans and reference or ground truth spans (in our case spans collected as a part of the editing process from writers). The overlap is measured at the character level, such that spans that partially overlap will get precision scores that reflect the amount of overlap between the two spans. High span-level precision indicates that the model is precise in identifying the correct boundaries of relevant text spans without over-predicting. We implement two precision metric variants: General and Categorical Precision. General Precision credits span selection regardless of category assignment, while Categorical Precision requires correct category assignment. We use a precision-based metric (like BLEU (Papineni et al., 2002)) rather than recall-based (like ROUGE (Lin, 2004)) as LLM-based methods tend to over-generate spans, which recall doesn’t penalize. Our focus is measuring overlap between generated and ground truth spans. Appendix A.2 provides a simplified example where General and Categorical Precision are computed, illustrating the suitability of these metrics for this evaluation setting.
General Precision Categorical Precision
Expert Agreement 0.57 0.23
Detector LLM 2-shot 5-shot 25-shot 2-shot 5-shot 25-shot
Claude3.5-Sonnet 0.43 0.46 0.44 0.20 0.21 0.20
Llama3.1-70b 0.42 0.45 0.38 0.16 0.21 0.14
GPT-4o 0.44 0.46 0.46 0.17 0.18 0.17
Table 9. Results of automated methods for detecting problematic spans in LLM-generated text, compared to agreement levels between three experts. Results report Precision scores for various LLMs used for detection, using a 2,5,25 examples in the few-shot prompt instruction, reporting both a General Precision and the stricter Categorical Precision.
We implement few-shot LLM-based methods (Brown et al., 2020) that have demonstrated competitive performance on tasks across several disciplines, often using fewer than 100 examples  [^9] Model fine-tuning experiments on LAMP Corpus are left for future work. Our experiment varies few-shot examples (2, 5, and 25) with the 2-shot prompt in Appendix A.1 and tests Llama3.1-70B, GPT-4o, and Claude3.5-Sonnet. As part of the collection of the LAMP Corpus, 50 paragraphs were edited independently, and we computed General and Categorical Precision on this set to estimate expert-expert agreement. Table 9 summarizes results. The best General Precision (0.46) is achieved by Claude-3.5 Sonnet and GPT-4o with a 5-shot prompt, below the expert agreement level (0.57). LLM-based methods can identify problematic spans with significant expert overlap, but improvement is possible. Performance improves from 2-shot to 5-shot prompts but plateaus thereafter (Claude-3.5 Sonnet and GPT-4o achieve similar or better performance with 5-shot vs. 25-shot prompts).
Categorical Precision is consistently lower than General Precision for both LLM-based methods and writer-writer comparison, suggesting that even when problematic spans are commonly identified, category agreement may differ. Section 5.2 explores this finding in detail, with examples of spans that could be reasonably assigned multiple categories. Table 10 illustrates the contrast between automatic editing and writer-selected edits by showing a paragraph with problematic spans and categories identified by a writer versus an LLM.
Writer
[…]
apparent advancement that framed his daily life, Jackson felt an inexplicable disconnection.
$\textbf{\color[rgb]{1,0,0}His}_{\textsc{{\color[rgb]{0,0,1}Unnecessary/Redundant Exposition}}}]$ mind often wandered to an untouched box of old photographs at home,
depicting simpler times – wooden houses, dirt paths, faces etched with stories of a bygone era. Those images
spoke to him more than the polished chrome and glass now surrounding him.[ He couldn’t help but yearn for the
intangible warmth of community gatherings under open skies, unhurried conversations, and the tangible
reality of a slower pace. Here, in the heart of progress, he was a mere observer, detached from the legacy
[…]
LLM
Jackson leaned back in his office chair, staring out through the expansive glass windows of the high-rise building.
[…]
$[\textbf{\color[rgb]{1,0,0}tangible reality of a slower pace}_{\textsc{{\color[rgb]{0,0,1}Purple Prose)}}}]$ Here, in the $[\textbf{\color[rgb]{1,0,0}heart of progress}_{\textsc{{\color[rgb]{0,0,1}Cliche}}}]$, he was a mere observer,
detached from the legacy beneath the steel and concrete, his own inclinations buried beneath the layers of
[…]
Table 10. Example of problematic spans identified by a writer and by a LLM.
6.2. Automatic rewriting of problematic spans in any LLM-generated text
Writer
Detected
and LLM
Rewritten
Dust settled on my window sill, a quiet testament to the daily upheaval outside a reminder of the
ongoing transformation outside. Yellow machines trundled along the narrow street, their relentless
growl a constant soundtrack growling incessantly. I watched the workers in their neon vests, like
bright insects against the concrete hive bright and busy against the concrete backdrop, orchestrating
a ballet of drills and hammers. Every so often, a stone house crumbled mansion fell to give way to sleek
modernism, steel skeletons reaching shyly towards a haze-obscured sky. The trees lining the avenue,
long sentries guarding our history, stood fewer each day The trees lining the avenue dwindled
each day. Neighbors nodded at one another with tight smiles exchanged tight smiles, masking the void
of vanished gardens and familiar creeks left by vanished gardens. Change buzzed in the air, metallic
and cold, seeping into conversations and dreams. Change felt cold and metallic, entering
conversations and dreams. I found myself wondering if we were wondered if we were building towards
something or away from itsomething or abandoning it.
LLM
Detected
and LLM
Rewritten
Becky’s mind races, her thoughts a tangled mess of worries and what-ifs. Becky’s thoughts spiral.
She knows she should eat, but the mere thought of food turns her stomach. she has no appetite. Instead,
she finds herself pacing the worn carpet of her small apartment, her fingers absently tracing the edges
of picture frames and trinkets. paces her small apartment, fingers tracing picture frames and trinkets
as she walks. She pauses at the window, watching the world outside continue its relentless march forward
go about its day. A neighbor walks their dog, oblivious to her inner turmoil oblivious. Becky’s gaze drifts
to the potted plants on her windowsill, their leaves drooping slightly. She reaches for the watering can,
tending to their needs as a way to quiet her ownnurturing them to distract herself.The simple act of
nurturing something else helps ground her, if only momentarily. The simple act grounds her
momentarily. She moves to her bookshelf, running her fingertips along the spines of well-loved novels,
searching for a familiar story to lose herself in hoping to find comfort in a familiar story. As she settles
into her favorite chair, book in hand, she notices her heartbeat has slowed. The knot in her stomach loosens,
ever so slightly. She exhales slowly. She may not be ready to eat yet, but she’s found a moment of peace in
the chaos of her mind. she’s found a small respite.
Table 11. Table showing LLM-edited paragraphs on both oracle Writer predicted spans as well as LLM predicted spans
To propose improvements for detected problematic spans, we use few-shot prompting (Brown et al., 2020) with LLMs as well. We design prompts for each of the seven edit categories, incorporating examples of rewrites from writers. Each prompt includes a category definition, 25 examples from the LAMP Corpus with original paragraphs, a single problematic span for the category, the expert-proposed rewrite, and finally the input paragraph with the target span to be rewritten. The prompts for each category are listed in Appendix A.3.
The detection and rewriting methods can form a two-step pipeline for editing paragraphs. Detection identifies problematic spans and assigns categories while rewriting uses category-specific prompts to revise each detected span. A final step replaces all problematic spans in the original paragraph with their rewrites. Unlike the detection task, we do not evaluate the rewriting stage in isolation. Instead, we judge the complete pipeline that edits an entire paragraph (by detecting and rewriting multiple spans) through manual evaluation with 12 writers that annotated the LAMP Corpus. We describe this manual experiment next.
6.3. Evaluating Automatic Editing of LLM-generated Text
To evaluate editing quality, we design an evaluation task where participants read three variants of a paragraph and rank them in terms of overall preference: (1) an unedited LLM-generated paragraph from the LAMP Corpus, (2) the Writer-edited version from the LAMP Corpus, and (3) an LLM-edited version using our pipeline to detect and rewrite problematic spans. We re-hired 12 of the 18 experts who had participated in creating the LAMP Corpus for this evaluation.
We split the LLM-edited variant into two further sub-conditions:
  • Writer Detected and LLM Rewritten: In this condition, the pipeline skips automatic detection of problematic spans, relying only on reference spans selected by the writer during manual editing. It runs solely the rewriting stage, simulating an oracle setting where problematic spans are manually provided. This condition is coded as LLM-edited-Oracle
  • LLM Detected and LLM Rewritten: In this condition, the two-step pipeline is entirely automatic, with the automatically detected spans being provided to the automatic rewriting module. This condition fully automates editing of the paragraph and is coded as LLM-edited-full.
Figure 7. Interface used by participants to read through variants of a paragraph (one LLM-generated, one manually edited by an expert, one edited by an LLM-based system), and rank them in terms of preference.
Table 11 shows examples of LLM-edited paragraphs under both sub-conditions. In our pilot evaluation, we initially included all four conditions for annotation. However, ranking four paragraphs proved challenging for participants, especially when distinguishing between their second and third preferences. Based on this feedback, we redesigned the task to have participants judge only three conditions in each annotation. We always included the LLM-generated and Writer-edited paragraphs and alternated between including LLM-edited-oracle and LLM-edited-full paragraphs. We note that to obtain automatic edits of a paragraph, we used the same LLM that had originally been used to generate the paragraph. [^10] In other words, we used GPT-4o in the two-step pipeline to generate edits to paragraphs that were originally generated by GPT-4o. While not optimal, as a single LLM might offer slightly better detection and rewriting capabilities, this approach allows us to simplify the experiment conceptually and and also test our hypothesis if edits lead to overall better alignment without relying on a single model family. We assess if using an LLM in a multi-stage pipeline (drafting, problem detection, rewriting) can enhance overall writing quality. Future work could potentially optimize this editing pipeline further, possibly yielding better results for LLM-edited conditions.
To ensure fairness, paragraph variants are displayed in a shuffled order and anonymized, and participants were not informed about the difference between the paragraphs (i.e. whether they are edited). For the curious reader, Figure 7 provides the interface used for the annotation task, including three variants of a paragraph. To conduct our experiments, we selected a total of 200 paragraph triplets (100 including an LLM-edited-oracle paragraph, and 100 including an LLM-edited-full paragraph) selecting samples from the LAMP Corpus’s test set. Preference judgments were collected in batches of 25-35 paragraph triplets, with participants paid \$35/hour. To account for potential subjectivity and calculate agreement and reliability, three experts judged each triplet, totaling 600 annotated preference rankings. To ensure the validity of the results, no participant reviewed paragraphs they had seen or edited in past tasks, and only judged paragraphs edited by other experts.
LLM-generated Writer-edited LLM-edited-full
2.55 1.47 1.99
LLM-generated Writer-edited LLM-edited-oracle
2.47 1.53 1.99
Table 12. Average Ranking across 600 preference judgments. LLM-edited $>$ LLM-generated (p-value: 1.3e-11 for Writer Predicted spans; 2.8e-13 for LLM Predicted spans) and Writer-edited $>$ LLM-generated (p-value: 1.1e-26 for Writer Predicted spans; 1.17e-31 for LLM Predicted spans) using Wilcoxon signed-rank test
Figure 8. Distribution of rankings for each variant in the preference annotation study. Annotators read three variants of a paragraph (Writer-edited, LLM-generated, and either LLM-edited-oracle or LLM-edited-full) and ranked them by preference (1st, 2nd, 3rd). The distribution indicates how often each variant was ranked as best (1st), second best (2nd), or worst (3rd).
To analyze the reliability of the results we calculate inter-annotator agreement using Kendall’s W (also known as Kendall’s coefficient of concordance) (Field, 2005) which ranges from 0 (no agreement) to 1 (complete agreement) to evaluate agreement amongst participants. Our annotation achieves an overall agreement of 0.505, suggesting a moderate level of agreement across all participants. This moderate agreement underscores the subjective nature of judging writing quality while suggesting that certain differences are distinctive enough to be consistently preferred by multiple participants.
Table 12 and Figure 8 summarize the preference evaluation results, showing average ranks across 600 annotations. Overall, the Writer-edited condition is most preferred, a sign that expert-edited text is unrivaled in terms of writing quality, being marked as the most-preferred paragraph variant 65% of the time and achieving an average rank of 1.5. Next, the LLM-edited variants come in second, with an average rank of 1.99 for both the LLM-edited-oracle and LLM-edited-full conditions. Surprisingly, the condition that leveraged the oracle span from writers ranks almost identical to the condition with automatically detected spans. This provides evidence that detection of problematic spans is not the bottleneck in improving writing quality, and instead the rewriting module (which is common to both conditions) is what dictates the overall performance of an automated text-editing pipeline. Finally, the original LLM-generated paragraphs achieve the worst ranking performance, being least preferred 60% of the time, and achieving an average rank of 2.51.
In summary, our experiment validates the potential benefit of automatic editing to improve writing quality: although automatic editing does match the quality of edits provided by professional writers, LLM-edited text is significantly preferred to LLM-generated text by expert writers (Design Principle 3). In other words, this experiment shows that LLMs can improve the quality of their writing in a fully automatic way, by first generating a draft, selecting problematic spans, and then rewriting such spans.
7. DISCUSSION
7.1. How is editing human writing different from LLM-generated text?
Editing human writing and LLM-generated text presents distinct challenges and requires different approaches. Human writing often contains nuanced expressions, personal style, and contextual references that reflect the author’s unique voice and experiences. Editors must preserve these elements while refining clarity, structure, and coherence. In contrast, LLM-generated text may lack consistent tone and exhibit repetitive patterns (Section 5). We asked writers to explain the differences in editing LLM-generated text compared to Human-written text. Several writers mentioned that LLM writing often required more extensive editing, mainly to remove unusual and sometimes nonsensical metaphors, inappropriate use of complex vocabulary that doesn’t fit the context and improving an overall tone that comes across as impersonal and mechanical. In a exchange W3 noted “I edit a lot of prose for my magazine but one thing that stuck with me as I was editing these paragraphs are the massive amount of cliché, histrionic descriptions, and direct exposition of intended meanings rather than effective representation. Indeed very strange”. They observed that the types of edits needed for LLM-generated text were often similar, but the sheer volume of necessary changes was higher than that with respect to human writing. LLM-generated content’s repetitive nature paradoxically made the human-editors feel robotic while trying to improve it.
7.2. How well can LLMs mimic edits from writers?
Our preference ranking results in Section 6.3 indicate that automatically edited paragraphs frequently rank second and sometimes first. This raises questions about LLMs’ ability to analyze textual patterns and generate content closely resembling a given writer’s edit. For the span “Janet lay in bed each night, her mind a whirlpool of restless thoughts”, both LLM and writer identified it as cliché. The LLM edited it to “Janet lay in bed each night, unable to sleep” while the writer changed it to “Each night, Janet lay prone in her bed and unable to sleep”. LLMs can also split run-on sentences and improve poor structure. For “Sarah froze, realizing it was her high school sweetheart, Alex, whom she hadn’t seen in over a decade”, the LLM edited it to “Sarah froze. It was Alex, her high school sweetheart. She hadn’t seen him in over a decade”, similar to the writer’s edit. However, LLMs sometimes replace clichés with other clichés or fail to remove unnecessary exposition. The most challenging edit category is Lack of Specificity and Detail where LLMs often fail to add engaging details. For “Her irritation slowly morphed into a strange, disconnected calm”, the writer added “After all, the noise just meant that she wasn’t the only one awake at this hour.” The model’s edit was less effective: “Her irritation slowly morphed into a strange, disconnected calm. The repetitive thump-thump-thump became almost hypnotic, lulling her into a trance-like state”. One potential limitation in our experiments is our reliance on few-shot instructions, requiring the model to learn rewriting from only a few examples. Training on the entire LAMP Corpus or more data might improve edit quality.
7.3. What recommendations can we provide for future LLM-based writing support tools that aspire to improve co-writing experience?
Eminent author Curtis Sittenfeld calls LLM writing the literary equivalent of fat-free cookies  [^11] https://www.nytimes.com/2024/08/20/opinion/beach-read-ai.html. LLMs are proficient at producing sentences that are grammatically correct and devoid of spelling errors. Beyond that, LLMs require extensive learning to effectively assist humans in improving their writing. In his essay Politics and English Language (Orwell, 2013), George Orwell said “Never use a long word when a short one will do.” LLM writing transgresses this simple rule by overusing lofty words. Clichés are bound to slip into even the best human writing, but when it comes to LLMs it simply cannot write without them. We believe this is partially a drawback of the technology behind LLMs. When a LLM calculates the probability of one word following another, clichés become very likely, because they’ve appeared so many times before. This explains why every other generated response is rife with clichés despite our prompt explicitly asking LLMs to avoid clichés and overused tropes (Table 13). LLMs need to learn how to identify and write without clichés such that it is engaging to every single reader. Overwriting is a bigger problem than underwriting. The rule for most writers is, “If in doubt, cut it.” (to Edit Your Own Writing, 2024) The Pulitzer Prize-winning writer John McPhee has called the process “writing by omission.” (McPhee, 2015). To become a better writer LLMs need to learn how to avoid unnecessary exposition. Last but not the least structure is what good writing hangs on (to Edit Your Own Writing, 2024). Long, run-on sentences are hard to read, and LLMs need to know when and how to split effectively to better manage flow and clarity.
7.4. What are the potential long-term effects on language evolution and writing styles as LLM becomes more prevalent and how can aligned editing tools help?
The increasing prevalence of large language models (LLMs) could significantly impact language evolution and writing styles over time. There’s potential for more homogenized writing as people rely on LLM-generated content, possibly leading to a reduction in linguistic diversity and individual voice (Shumailov et al., 2024). However, well-designed editing tools aligned with expert writing practices could help counteract these effects. Such tools could encourage more nuanced and sophisticated language use, preserve stylistic diversity, and promote critical thinking about word choice and sentence structure. By highlighting elements of expert writing, these tools could elevate overall writing quality while still allowing for personal expression, potentially steering language evolution towards greater clarity, precision, and effectiveness in communication (Li et al., 2024).
8. Limitations
While our study provides valuable insights into improving LLM-generated text through expert editing, there are several limitations to consider. Our study was conducted with 18 MFA-trained creative writers. While this ensured a high level of expertise, it may limit the generalizability of our findings. Future research could expand the participant pool to encompass diverse cultural backgrounds and writing traditions. The editing data primarily comes from literary fiction and creative non-fiction, making the identified idiosyncrasies and editing strategies potentially less applicable to other genres like technical writing, journalism, or scientific writing. Future work can expand on the line of work by including a broader range of writing styles and purposes. The selected LLMs (GPT-4, Claude 3.5 Sonnet, and Llama 3.1) are among the most advanced models, but they might not fully represent the entire spectrum of AI writing abilities. It should be noted that the evaluation of writing quality is inherently subjective, even with multiple annotators and inter-annotator agreement calculations. Experts may disagree on what constitutes an improvement, potentially influencing our results and their interpretation. Our automated methods for detecting and rewriting problematic spans relied on few-shot learning with a limited number of examples. While this approach showed promise, it may not fully capture the complexity and nuance of expert editing and training a model on the entire LAMP Corpus or additional data is required.
It should be noted that while paragraph-level editing provides a balance between granularity and context, it may miss broader structural or thematic issues that become apparent only when considering longer pieces of writing. Last but not least, our study relied on professional writers editing AI-generated text for monetary compensation, which may have influenced the quality and nature of the edits. Editing one’s own work typically involves more personal investment than editing text for pay, potentially leading to less motivation for substantial improvements (Byron and Khazanchi, 2012). Additionally, the repetitive nature of editing multiple AI-generated paragraphs could lead to fatigue, especially if the content is perceived as uninteresting or lacking in creativity. This fatigue could result in less thorough or thoughtful edits as the task progresses.
9. Conclusion
In this work, we present a comprehensive approach to mitigating idiosyncrasies and improving human-AI alignment in the writing process through expert editing. We i) develop a taxonomy of edit categories grounded in established writing practices, ii) create the LAMP corpus containing over 8,000 fine-grained edits by professional writers on LLM-generated text, and iii) design methods for automatic detection and rewriting of problematic spans. Our analysis reveals several key findings. Professional writers identify consistent categories of edits needed to improve AI writing. Surprisingly, there are no significant differences in perceived writing quality or types of edits needed across texts generated by different large language models (GPT-4, Claude 3.5, Llama 3.1). Automated methods using few-shot prompting are able to detect and rewrite problematic spans in LLM generated text, though far from matching human expert performance. Finally, in terms of preference evaluations, writers consistently rank text edited by other writers highest, followed by LLM-edited text, with unedited LLM-generated text ranking lowest. As AI text generation becomes more prevalent, developing robust editing and alignment techniques will be crucial to ensure AI systems produce high-quality writing that meets human standards and enhances creativity and linguistic diversity.
Appendix A Appendix
Figure 9. Interface for formative study to collect fine-grained labels
[…]
Response
Prompt
Imagine you are a fiction writer for the NewYorker. Now write a paragraph (10-15 sentence)
as a response to the following question. Try your best to be original, avoiding clichés or
overused tropes. Do not use ornamental language and focus on nuance, simplicity, and subtext.
[…]
Imagine you are a writer for the New York Times Modern Love section. Now write a
paragraph (10-15 sentence) as a response to the following question. Try your best to be
original, avoiding clichés or overused tropes. Do not use ornamental language and focus
[…]
Imagine you are a writer for the New York Times Cooking section. Now write a paragraph
(10-15 sentence) as a response to the following question. Try your best to be original, avoiding
clichés or overused tropes. Do not use ornamental language and focus on nuance, simplicity,
[…]
Imagine you are a writer for the New York Times Travel section. Now write a paragraph
(10-15 sentence) as a response to the following question. Try your best to be original, avoiding
clichés or overused tropes. Do not use ornamental language and focus on nuance, simplicity,
[…]
Imagine you are a beloved female Internet advice columnist whose trademark is deeply felt
and frank responses grounded in your own personal experience. Now write a paragraph
(10-15 sentence) as a response to the following question.Try your best to be original, avoiding
clichés or overused tropes. Do not use ornamental language and focus on nuance,
[…]
Table 13. Prompts for generating instructions and responses
A.1. Idiosyncracy Span Detection Prompt
{mdframed}
[ linecolor=black, linewidth=1pt, innertopmargin=6pt, innerbottommargin=6pt, innerrightmargin=10pt, innerleftmargin=10pt, frametitle=Idiosyncracy Span Detection Prompt, frametitlebackgroundcolor=black, frametitlefont=, ] You are given a paragraph of writing, and your goal is to provide feedback by selecting spans of text in the writing that could be improved, and assign each problematic span to an error category. Below, we list the 7 error categories that you can choose from.
You are also provided 2 examples of paragraphs that were annotated by professional writers, which you can use to better understand the task and the error categories.
Error Categories:
- ”Awkward Word Choice and Phrasing”: Suggestions for better word choices or more precise phrasing to enhance clarity and readability.
- ”Cliche”: The use of hackneyed phrases or overly common imagery that lack originality or depth.
- ”Poor Sentence Structure”: Feedback on the construction of sentences, recommending changes for better flow, clarity, or impact.
- ”Unnecessary/Redundant Exposition”: Redundant or non-essential parts of the text that could be removed/rephrased for conciseness.
- ”Lack of Specificity and Detail”: Need for more concrete details or specific information to enrich the text and make it more engaging.
- ”Purple Prose”: Identifying parts of the text that are seen as unnecessary ornamental and overly verbose.
- ”Tense Consistency”: Comments pointing out inconsistencies in verb tense that need to be addressed for uniformity.
Example 1: Input Text
Output:
Example Output in JSON format.
Example 2:
(Similar to example 1)
Rules:
- Number of Spans – You can provide feedback on multiple spans, and multiple spans can have the same category.
- Span must be verbatim – The span you select must be verbatim from the paragraph, otherwise, the feedback will not be provided to the user.
- No Overlap – Spans should not overlap, and one span should not include the other.
- Single Category – Each span should have exactly one category from the categories listed above.
Paragraph:
PARAGRAPH
A.2. Precision Metrics Explanation and Example
We illustrate the General and Categorical Precision on a simple example and justify the choice of the metric.
Imagine we have the following sentence, that has been annotated by a human annotator: On this dark and stormy night, her heart skipped a beat as she was afraid of what was to come.
ANNOTATION = On this dark and stormy night[CLICHÉ], her heart skipped a beat as she was afraid of what was to come.[UNNECESSARY EXPOSITION]
Now let’s imagine that System 1 and System 2 have produced the following predictions:
SYSTEM 1 = On this dark and stormy night, her heart skipped a beat as she was afraid of what was to come.[CLICHÉ]
SYSTEM 2 = On this dark and stormy night[CLICHÉ], her heart skipped a beat as she was afraid of what was to come.[CLICHÉ]
We extract the annotated spans:
  • Span 1: characters [9,30]; category: CLICHÉ
  • Span 2: characters [57, 94]; category: UNNECESSARY EXPOSITION
System 1 produced a single span:
  • Span 1: characters [9, 94]; category: CLICHÉ
System 2 produced two spans:
  • Span 1: characters [19,30]; category: CLICHÉ
  • Span 2: characters [57, 94]; category: CLICHÉ
We can first compute General Precision, which disregards the category of the spans. It is the overlap between predicted spans and annotated spans, divided by the total amount of predicted characters:
  • General Precision (System 1) = ((30-9) + (94-57)) / (94-9) = 0.68
  • General Precision (System 2) = ((30-19) + (94-57)) / ((30-19) + (94-57)) = 1.0
System 2 achieves a higher precision, as all the spans it predicted were included in the manual annotation. On the other hand, System 1 predicted a larger span that included the annotated span, but also additional characters, causing a lower precision score.
When consider Categorical Precision, overlap is only considered as valid if the overlapping spans coincide in category. The scores would be:
  • Categorical Precision (System 1) = (1*(30-9) + 0*(94-57)) / (94-9) = 0.25
  • Categorical Precision (System 2) = (1*(30-19) + 0*(94-57)) / ((30-19) + (94-57)) = 0.23
System 1 achieved higher categorical precision by fully overlapping with the annotated CLICHÉ span, while System 2 only partially overlapped. Both systems incorrectly categorized the second span, resulting in lower precision scores.Precision scores can be inflated by reducing predictions, but our LLMs weren’t instructed to optimize for precision. In fact, they tend to select more spans than human annotators, leading to high recall but potentially lower precision. We focus on precision to penalize systems that produce too many or overly large spans.
A.3. Rewriting Prompts
A cliché is a saying, idea, or element of an artistic work that has become overused to the point of losing its
original meaning or effect, even to the point of being weird, irritating, or bland
You will be given example of 25 paragraphs with spans that count as Cliche and suggested edits that either
*REWRITES THE CLICHE or SIMPLY REMOVES IT**.
Your task will then be to suggest edits (either spans or empty string) that gets rid of the cliche while making
the resulting paragraph coherent, given a new paragraph and highlighted span of Cliche from it. Do not simply
paraphrase or use fancy ornamental language; Try to keep each sentence short. Look at the examples carefully
*IT IS VERY IMPORTANT TO MAKE SURE THAT YOUR EDITED TEXT ONCE ADDED TO THE
PARAGRAPH READS COHERENTLY AND GRAMMATICALLY CORRECT.
For instance if you replace text within <span></span>tags with a longer span; please make sure the following
text after the edit, is its continuation. Simple way to ensure this is to make sure that the edited span has the same
casing and punctuation at the beginning and end as that of the original span.
PLEASE FOLLOW THE OUTPUT SCHEMA AS THE EXAMPLES BELOW AND DO NOT RETURN
ANYTHING OTHER THAN THE EDITED SPAN WITHIN QUOTES
Example 1
Paragraph: Matthews had lived in the Valley all his life, and its rhythms and secrets were etched into his
being <span>like the lines on a well-worn map</span>. He knew […]
Original Span: ”like the lines on a well-worn map”
Edited Span: ”like creases in an old pocket map”
Example 18
Paragraph: Husna sat at the ancient wooden [….] <span>The room was a bubble of quiet concentration, the
only sounds the clacking of the typewriter, the rustling of paper, and the occasional whistle of the teakettle in
the adjoining kitchen.</span>
Original Span: ”The room was a bubble of quiet concentration, the only sounds the clacking of the typewriter,
the rustling of paper, and the occasional whistle of the teakettle in the adjoining kitchen.”
Edited Span: ”The room was quiet. The outside world did not exist. At times, Husna tapped her foot. Shah
Sahib coughed and she would stop. The typewriter never did.”
Example 25
Paragraph: Last night, I dreamt of an [….] She didn’t speak, but her eyes <span>communicated a haunting mix
of sadness and knowing, as if she held</span>the weight of forgotten secrets. I felt a […]
Original Span: ”communicated a haunting mix of sadness and knowing, as if she held”
Edited Span: ”conveyed”
Table 14. Prompt to rewrite Cliche
Poor sentence structure refers to writing that is difficult to understand or lacks clarity due to issues with how
sentences are constructed. It encompasses issues like run-on sentences, fragments, misplaced or dangling
modifiers, lack of variety, overuse of passive voice, improper parallelism, and unclear pronoun references,
all of which impede clear communication and reader comprehension
You will be given example of 25 paragraphs with text within <span></span>tags that shows poor sentence
structure and suggested edits that either **REWRITES WITH IMPROVED SENTENCE STRUCTURE**.
Your task will then be to suggest edits that rewrites the text within the span tags with better sentence structure
while making the resulting paragraph coherent, given a new paragraph and highlighted span of poor sentence
structure from it. Do not use fancy ornamental language; Look at the examples carefully and do not output
anything after closing quotes.
*IT IS VERY IMPORTANT TO MAKE SURE THAT YOUR EDITED TEXT ONCE ADDED TO THE
PARAGRAPH READS COHERENTLY AND GRAMMATICALLY CORRECT. For instance if you replace
text within <span></span>tags with a longer span; please make sure the following text after the edit, is its
continuation.
PLEASE FOLLOW THE OUTPUT SCHEMA AS THE EXAMPLES BELOW AND DO NOT RETURN
ANYTHING OTHER THAN THE EDITED SPAN WITHIN QUOTES
Example 4
Paragraph: <span>As the night wore on, Z.’s laughter grew louder, his words slurring together like a sloppy
melody. N. and I exchanged a knowing glance, our concern simmering beneath the surface.</span>At first, it
was just a slight stumble, a misstep that could be brushed off as a joke. […..]
Original Span: ”As the night wore on, Z.’s laughter grew louder, his words slurring together like a sloppy melody.
N. and I exchanged a knowing glance, our concern simmering beneath the surface.”
Edited Span: ”Z. was drinking more and more as the night went on. He laughed more loudly. His words started to
slur, blurring one into the next. I looked at N., who knew what I was thinking. We were going to have to take
care of him.”.
Example 13
Paragraph: <span>As I step into the quiet, garden-facing room on the second floor, I’m struck by the sense of stillness
that pervades the space</span>. The occupants, an elderly couple, sit motionless in their armchairs, their [….]
Original Span: ”As I step into the quiet, garden-facing room on the second floor, I’m struck by the sense of
stillness that pervades the space”
Edited Span: ”A sense of stillness pervades the garden-facing room on the second floor” .
Example 25
Paragraph: Chef Amelia raced […..] <span>She plastered on a polite smile, determined not to let her personal history
interfere with her professional duties.</span>As Daniel approached, plate in hand, Amelia steeled herself […..]
Original Span: ”She plastered on a polite smile, determined not to let her personal history interfere with her
professional duties.”
Edited Span: ”She shot a dutiful smile for anyone who was looking. This was an important night, and she wasn’t
going to let the past get in the way of a job well done.”
Table 15. Prompt to rewrite Poor Sentence Structure
Unnecessary or redundant exposition in writing refers to providing excessive explanatory information that
doesn’t contribute meaningfully to the story, characters, or overall narrative.
You will be given example of 25 paragraphs with text within <span></span>tags that count as
unnecessary/redundant exposition and suggested edits that either **REWRITES IT IN FEWER WORDS
or SIMPLY REMOVES IT**.
Your task will then be to suggest edits that rewrites the text within the span tags correcting the unnecessary
/redundant exposition while making the resulting paragraph coherent, given a new paragraph and
highlighted text within of unnecessary/redundant exposition. Do not simply paraphrase or use fancy
ornamental language or repeat the same thing in the edited span; Look at the examples carefully.
*IT IS VERY IMPORTANT TO MAKE SURE THAT YOUR EDITED TEXT ONCE ADDED TO THE
PARAGRAPH READS COHERENTLY AND GRAMMATICALLY CORRECT.
For instance if you replace text within <span></span>tags with a shorter span; please make sure the
following text after the edit, is its continuation. Simple way to ensure this is to make sure that the edited
span has the same casing and/or punctuation at the beginning and end as that of the original span.
PLEASE FOLLOW THE OUTPUT SCHEMA AS THE EXAMPLES BELOW AND DO NOT RETURN
ANYTHING OTHER THAN THE EDITED SPAN WITHIN QUOTES
Example 2
Paragraph: In spring, when the first buds unfurled […] embrace of varenyky dinners provided comfort against
the chill <span>, each bite narrating a history of resilience and hope</span>. It was through […]
Original Span: ”, each bite narrating a history of resilience and hope”
Edited Span: ””
Example 18
Paragraph: <span>As Oghi watched his mother-in-law, Mrs. Kim, he felt a subtle sense of unease settle in the
pit of his stomach.</span>It wasn’t just the uncharacteristic behavior itself - […]
Original Span: ”As Oghi watched his mother-in-law, Mrs. Kim, he felt a subtle sense of unease settle in the pit
of his stomach.”
Edited Span: ”Oghi watched his mother-in-law Mrs. Kim with heightening unease.”
Example 23
Paragraph: The small room [….] They teased and corrected each other’s recollections <span>, creating a tapestry
of resilience and camaraderie</span>.It wasn’t all smooth-sharp words resurfaced around old wound, [….]
Original Span: ”, creating a tapestry of resilience and camaraderie”
Edited Span: ””
Table 16. Prompt to rewrite Unnecessary or redundant exposition
Lack of Specificity and Detail in writing refers to the absence of concrete and specific information, which can make the text
feel vague and unengaging. The need for more concrete details or specific information is crucial to enrich the text and make
it more engaging. Specificity helps to create vivid imagery, provides clarity, and connects with the reader on a deeper level.
doesn’t contribute meaningfully to the story, characters, or overall narrative.
You will be given example of 25 paragraphs with text within <span></span>tags that lacks specificity and detail and suggested
edits that either **REWRITES WITH SPECIFICITY AND DETAIL**.
Your task will then be to suggest edits that rewrites the text within the span tags with specificity and detail that is engaging while
making the resulting paragraph coherent, given a new paragraph and highlighted span of lack of specificity and detail from it.
Do not simply paraphrase or use fancy ornamental language; Look at the examples carefully and do not output anything after
closing quotes.
*IT IS VERY IMPORTANT TO MAKE SURE THAT YOUR EDITED TEXT ONCE ADDED TO THE PARAGRAPH READS
COHERENTLY AND GRAMMATICALLY CORRECT. For instance if you replace text within <span></span>tags with a longer
span; please make sure the following text after the edit, is its continuation. Simple way to ensure this is to make sure that the edited
span has the same casing and punctuation at the beginning and end as that of the original span.
PLEASE FOLLOW THE OUTPUT SCHEMA AS THE EXAMPLES BELOW AND DO NOT RETURN
ANYTHING OTHER THAN THE EDITED SPAN WITHIN QUOTES
Example 1
Paragraph: Sarah Mitchum’s marriage appeared outwardly conventional, but subtle tensions simmered beneath the surface. She
and […..] leaving Sarah feeling increasingly isolated within her <span>own marriage.</span>
Original Span: ”within her own marriage.”
Edited Span: ”.Their marriage had run its course. There was no coming back.”
Example 15
Paragraph: <span>Dr. Arthur Steiger’s fall from grace began with a series of whispered concerns among his colleagues at
Cormac General Hospital.</span>The small-town pain specialist had always been known [….]
Original Span: ”Dr. Arthur Steiger’s fall from grace began with a series of whispered concerns among his colleagues at Cormac
General Hospital.”
Edited Span: ”Pain was Dr. Arthur Steiger’s forte. Not inflicting it, that is, but resolving it. Whenever a patient had problem,
whether a tear in atendon, a sprain, a knock, a headache, a broken bone– it was Dr. Steiger that knew what to do.”
Example 21
Paragraph: Mila sat on her porch a week after the storm had hit, sipping lukewarm tea. [….] Each night <span>it grew louder,
shifting from a whisper to a groan, but she had dismissed it, too tired from long days at work</span>. […..]
Original Span: ”it grew louder, shifting from a whisper to a groan, but she had dismissed it, too tired from long days at work”
Edited Span: ”lying like blanched spinach in her IKEA bed, trying not to think about another day of writing emails with someone
else’s signature on them and pretending not to care what John Blanchett, CEO of Executive Industries thought of her blouse–in
other words,another day as John’s executive assistant–”
Table 17. Prompt to rewrite Lack of Specificity and Detail
In literary criticism, purple prose is overly ornate prose text that may disrupt a narrative flow by drawing undesirable attention
to its own extravagant style of writing, thereby diminishing the appreciation of the prose overall.Purple prose is characterized
by the excessive use of adjectives, adverbs, and metaphors.
You will be given example of 25 paragraphs with text within <span></span>tags that has purple prose in it and suggested edits
that either **REWRITES THEM WITH SIMPLER WORDS OR REMOVES IT**.
Your task will then be to suggest edits that rewrites the text within the span tags altering the purple prose while making the
resulting paragraph coherent, given a new paragraph and highlighted span of purple prose from it. Do not simply paraphrase
or use fancy ornamental language; Look at the examples carefully and do not output anything after closing quotes.
*IT IS VERY IMPORTANT TO MAKE SURE THAT YOUR EDITED TEXT ONCE ADDED TO THE PARAGRAPH
READS COHERENTLY AND GRAMMATICALLY CORRECT. For instance if you replace text within <span></span>
tags with a longer span; please make sure the following text after the edit, is its continuation. Simple way to ensure this is
to make sure that the edited span has the same casing and punctuation at the beginning and end as that of the original span.
PLEASE FOLLOW THE OUTPUT SCHEMA AS THE EXAMPLES BELOW AND DO NOT RETURN ANYTHING
OTHER THAN THE EDITED SPAN WITHIN QUOTES
Example 2
Paragraph: <span>Fruto never intended to stir anything beyond the melting pot of their weekly card game.</span>But
when the chatter turned to the dry monotony of their jobs, Fruto found himself blurting out, [….]
Original Span: ”Fruto never intended to stir anything beyond the melting pot of their weekly card game.”
Edited Span: ”Fruto hadn’t meant to disrupt the routine of their weekly card game.”
Example 16
Paragraph: My mother cried, [….] All of it vanished<span>, cycling back through her mind, not as numbers but memories
of scraped knees she bandaged alone and birthdays where her absence was felt more acutely than her presence. The sobs
emerged from this deep well of unspoken expectations, leaving behind a residue of weary resilience and a few hopeful
echoes yet unwilling to completely extinguish.</span>
Original Span: ”, cycling back through her mind, not as numbers but memories of scraped knees she bandaged alone and
birthdays where her absence was felt more acutely than her presence. The sobs emerged from this deep well of unspoken
expectations, leaving behind a residue of weary resilience and a few hopeful echoes yet unwilling to completely extinguish.”
Edited Span: ”She cried. She cried deep from this well of scraped knees she bandaged alone and birthdays she missed to
work. She cried for unfairness. She cried without relief.” .
Example 24
Paragraph: <span>As they navigated their final year of high school, Maya and Jake found themselves at a crossroads, their
educational paths diverging like tributaries of a river.</span>[….]
Original Span: ”As they navigated their final year of high school, Maya and Jake found themselves at a crossroads, their
educational paths diverging like tributaries of a river.”
Edited Span: ”The final year of high school was pulling Maya and Jake in different directions.”
Table 18. Prompt to rewrite Purple Prose